RSES/ph222/W1 HW.pdf WI An electric drill starts from rest and rotates with a constant angular acceleration. After the drill has rotated through a certain angle, the magnitude of the centripetal acceleration of a point on the drill is twice the magnitude of the tangential acceleration. What is the angle?
Answer #1

given

ac = 2at

where c is for centripetal and t is for tangential

ac = rw2

at = r\alpha

rw2 = 2r\alpha

w2 = 2\alpha

Using rotation kinematics, we have

2\alpha = 2\alpha*\Delta \theta

\Delta \theta = 1 rad